OpenStudy (anonymous):

If f(x) = -x+5 and g(x) = x^2, what is the solution set of the equation f(x)=g(x), rounded to the nearest tenth?

zepdrix (zepdrix):

\[\large\rm -x+5=f(x)=g(x)=x^2\]Leads us to,\[\large\rm -x+5=x^2\]Solve for x :)

OpenStudy (anonymous):

k

zepdrix (zepdrix):

Add x to each side, subtract 5,\[\large\rm 0=x^2+x-5\]Try to factor, if it doesn't factor, Quadratic Formula, ya? :)

OpenStudy (anonymous):

ok i'll factor ty

OpenStudy (anonymous):

i'll try completing the square

OpenStudy (anonymous):

(x+1/2)^2) -(21/4)

OpenStudy (anonymous):

oh noes

OpenStudy (anonymous):

u left :(

zepdrix (zepdrix):

oh sorry :O i get notification when u post though, so easy to come back

OpenStudy (anonymous):

Oh :)

zepdrix (zepdrix):

You did completing the square? Ok that's a fine approach,\[\large\rm 0=\left(x+\frac{1}{2}\right)^2-\frac{21}{4}\]Adding 21/4 to both sides,\[\large\rm \frac{21}{4}=\left(x+\frac{1}{2}\right)^2\]Then square root next, ya?

zepdrix (zepdrix):

Quadratic Formula would have been a bit easier than this route XD But we're already half-way through this hehe

OpenStudy (anonymous):

x+ 1/2 = sqrt 21/4

OpenStudy (anonymous):

but i can do quad form

OpenStudy (anonymous):

(-1 +or- sqrt 21)/2

zepdrix (zepdrix):

Ok great! And it looks like they want us to approximate our solutions to the nearest tenth. So, calculator time.

OpenStudy (anonymous):

yes

zepdrix (zepdrix):

\[\large\rm \frac{-1-\sqrt{21}}{2}\approx?\] \[\large\rm \frac{-1+\sqrt{21}}{2}\approx?\]

OpenStudy (anonymous):

1.79 2.79

zepdrix (zepdrix):

Hmm, one of those should be negative.

OpenStudy (anonymous):

-2.79

OpenStudy (anonymous):

yes?

zepdrix (zepdrix):

Ok great :) And make sure you round down to the correct place. Right now you have the `hundredth place`

OpenStudy (anonymous):

1.80 -2.80

zepdrix (zepdrix):

Don't include the 0, bad bad bad! XD That's still the hundredth's place hehe. 1.8 and -2.8 Yay good job \c:/

OpenStudy (anonymous):

oh i was worry

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

I appreciate

OpenStudy (anonymous):

I can become a fan because I already did ;p